OpenStudy (anonymous):

PLEASE HELP!! acceleration depends on time t therefore a=6t the body starts from rest .. find the average velocity between t=1 and t=5

OpenStudy (anonymous):

wrong answer .. sorry

OpenStudy (anonymous):

the correct answer is .... a=dv/dt=d2x/dt2 solving that you will get x=124 time=4 avg velocity =x/4=124/4 therefore answer equal to 31 m/s ... this is also the answer given by the book ..

OpenStudy (***[isuru]***):

OH! GOD AM SOOO SRY !!!

OpenStudy (***[isuru]***):

I HAVE miscalculate something!

OpenStudy (anonymous):

its ok man anyways im preparing for iit

OpenStudy (***[isuru]***):

OK!!! wish u best of luck!! for wtever u r preparing!

OpenStudy (dan815):